# Uniform irreducibility of Galois action on the $\ell$-primary part of Abelian $3$-folds of Picard type

### v1

@@ 1.4. Higher Iwahori invariants via matrix coefficients
In this subsection we assume that $E/\mathbf{Q}_{p}$ is unramified. The unique unramified Arthur packet is given by $\Pi(\lambda_{0})=\Pi(\lambda_{0},\mathbf{1})$, where $\lambda_{0}$ is the unique quadratic unramified character of $E^{\times}$.
It follows from Iwasawa decomposition that the corresponding $\operatorname{Ind}_{B}^{G}(\mu_{0})$ has one dimensional invariants by any given maximal open compact subgroup $K$ of $G$. The following proposition states, depending on $K$, whether the $K$-invariant line belongs to $\pi_{2}$ or maps non-trivially to $\pi_{n}$. We recall that $K^{\circ}$ and $K^{\prime}$ are the two standard maximal compact subgroups, $K^{\circ}$ being the hyperspecial one, and that the standard Iwahori subgroup $I$ equals $K^{\circ}\cap K^{\prime}$.
@@
One has $\dim\pi_{2}^{K^{\prime}}=\dim\pi_{n}^{K^{\circ}}=1$, if $\mu$ is unramified, and $\pi_{2}^{I}=\pi_{n}^{I}=\{0\}$, otherwise.
@@
Applying the Jacquet functor to the exact sequence (2) allows one to see that both $\pi_{2}$ and $\pi_{n}$ have non-trivial $I$-invariants if $\mu$ is unramified, and none, otherwise. Assuming henceforth that $\mu=\mu_{0}$ is unramified, we observe that both $\pi_{2}^{I}$ and $\pi_{n}^{I}$ must be $1$-dimensional. Moreover, as by Cartan decomposition $G$ is generated by $K$ and $\gamma$, hence by $K^{\circ}$ and $K^{\prime}$, it follows that necessarily one amongst $\pi_{2}^{I}$ and $\pi_{n}^{I}$ is fixed by $K^{\circ}$, while the other one is fixed by $K^{\prime}$.
By Iwasawa decomposition, the restriction of $\pi=\operatorname{Ind}_{B}^{G}(\mu_{0})$ to $K$ is given by $\operatorname{Ind}_{B\cap K}^{K}(\mu_{0})$, hence the line $\pi^{K}$ admits a basis $f_{K}$ characterized by asking its restriction to $K$ to be $\mathbf{1}_{K}$. Moreover, the line $(\pi^{\vee})^{K}$ admits a basis $f_{K}^{\vee}$ characterized by
$\langle f,f_{K}^{\vee}\rangle=\frac{1}{\sqrt{\operatorname{vol}(K)}}\int_{K}f(k)dk.$
@@
$\sqrt{\operatorname{vol}(K\gamma^{n}K)}\langle\gamma^{n}\cdot f_{K},f_{K}^{\vee}\rangle=\sqrt{[K:(K\cap\gamma^{-n}K\gamma^{n})]}\int_{K}f_{K}(k\gamma^{n})dk.$
Using Iwahori decomposition for all $n\geqslant 1$ we have $K:(K\cap\gamma^{-n}K\gamma^{n})]/[K:I]=p^{4n-3}$ (resp. $p^{4n-1}$), where $K=K^{\circ}$ (resp. $K^{\prime}$). Since $(\mu\delta^{1/2})(\gamma)=-p^{3}$ we have $f_{K}\in\pi_{2}$ if, and only if,
(5) $(\Phi^{K}_{n})_{n}\in\operatorname{L}^{2}(\mathbf{Z}_{\geqslant 0})\text{, where }\Phi^{K}_{n}=p^{5n}\cdot\int_{K}f_{K}(\gamma^{-n}k\gamma^{n})dk.$
The proof of Proposition is then completed by the following Lemma. ∎
@@
The last row of an element $k\in K^{\prime}$ is given by $(0,0,1)\cdot k=(p\cdot c_{1}(k),p\cdot c_{2}(k),c_{3}(k))$ with $c_{2}(k)\in\mathcal{O}$ and $(c_{1}(k),c_{3}(k))\in(\mathcal{O}\times\mathcal{O})\backslash(\mathcal{P}\times\mathcal{P})$. For $j\geqslant 0$ we let
[…]
@@
First, we compute the volume of $K^{\prime}_{j}$, for $j\geqslant 1$. Using Iwahori decomposition one finds that:
@@
Therefore $\frac{1}{\operatorname{vol}(K^{\prime})}\int_{K^{\prime}}f_{K}(\gamma^{-n}k\gamma^{n})dk=\frac{1}{[K^{\prime}:K^{\prime}_{2n}]}+p^{-6n}\sum_{j=0}^{2n-1}(-1)^{j}p^{3j}\frac{1}{[K^{\prime}:K^{\prime\times}_{j}]}=$
@@
Alternatively, one could use MacDonald’s formula for zonal spherical functions to see that $\pi_{2}$ does not admit non-zero vectors fixed by the hyperspecial maximal open compact subgroup $K^{\circ}$. Indeed, [15, §5.5] allows to express $\Phi^{K^{\circ}}_{n}$ from (5), up to a non-zero constant, as
$\Gamma_{\mu}\cdot\mu(\gamma^{-n})+\Gamma_{\mu^{w}}\cdot\mu^{w}(\gamma^{-n}),\text{ where }\Gamma_{\nu}=\frac{1-p^{2}\cdot\nu(\gamma^{-2})}{1-\nu(\gamma^{-2})}\cdot\frac{1-p^{2}\cdot\nu(\gamma^{-1})}{1-\nu(\gamma^{-1})},$
where the two factors in $\Gamma_{\nu}$ correspond respectively to the positive roots $\zeta$ and $2\zeta$ of $G$ (see§1.1). As $\mu(\gamma^{-1})=\mu^{w}(\gamma)=-p^{-1}$, one has $\Gamma_{\mu}=0\neq\Gamma_{\mu^{w}}$, hence the sequence $(\Phi^{K^{\circ}}_{n})_{n\geqslant 0}$ is not $\operatorname{L}^{2}$. This also shows, in passing, that $\pi_{n}$ is not square integrable.
@@
For $r\geqslant 1$ and for $\pi\in\Pi(\lambda_{0})$, one has $\dim\left(\pi^{I_{2r}}\right)\geqslant r+1$.
@@
By Proposition 1.6 for all $r\in\mathbf{Z}$, $\pi_{2}$ contains a (unique) line fixed by $\gamma^{r}K^{\prime}\gamma^{-r}$, having $\gamma^{r}\cdot f$ as basis, and moreover, the stabilizer in $G$ of that line is $\gamma^{r}K^{\prime}\gamma^{-r}$. We claim that the vectors $f,\gamma\cdot f,\dots,\gamma^{r}\cdot f\in\pi_{2}$ are linearly independent. Indeed, if $f$ was a linear combination of the remaining $r$ vectors then it would be fixed by $\cap_{1\leqslant j\leqslant r}\gamma^{i}K^{\prime}\gamma^{-i}$ of $G$ which is not contained in $K^{\prime}$. As any of these $(r+1)$ vectors is fixed by $I_{2r}$, the claim follows for $\pi_{2}$.
Arguing the exact same way, using $K^{\circ}$ instead of $K^{\prime}$, proves the statement for $\pi_{n}$. ∎
While using the unramified $A$-packet $\Pi(\lambda_{0})$ would be sufficient in our global applications when $D\equiv 3\pmod{8}$, the case of discriminants $D\equiv 7\pmod{8}$ would require the use of certain tamely ramified $A$-packets $\Pi(\lambda)$ and providing explicit levels for them is the object of the next subsection.
1.5. Intertwining and an exponential sum
Our arithmetic applications will require to show existence of non-zero $I_{r}$-invariants, for some $r\in\mathbf{Z}_{\geqslant 1}$, in certain ramified $A$-packets. This is delicate because of the lack of new-vector theory for non-generic representations (see Remark 1.12). Also Casselman’s result asserting that $\pi^{I_{r}}$ surjects onto $\pi_{N}^{T\cap I_{r}}$ is inconclusive here as the latter vanishes, contrarily to [10] where the open compact is a pro-$p$-Iwahori subgroup of a sufficiently deep level (see §4.3 where these results are discussed). We will instead resort to explicit methods to prove in Proposition 1.10 that $\pi_{n}$ has non-zero invariants $K_{T}$, which contains a conjugate of $I_{3}$. It is relatively straightforward to determine all such vectors $f$ in the full induced representation but it becomes a thorny issue to find a non-square integrable matrix coefficient $\langle g\cdot f,f^{\vee}\rangle$. By another result of Casselman, matrix coefficients can be expressed in terms of the corresponding ones in the Jacquet module, given here by an explicit scalar product on $\mathbf{C}\cdot\mu\oplus\mathbf{C}\cdot\mu^{w}$. Making this actually work requires non-vanishing under the Jacquet functor which, once verified, leads directly to the result we seek. The computation of the Jacquet functor is first reduced to a precise statement about the intertwining operator at the level of finite reductive groups. It involves showing non-vanishing of some explicit exponential sums, bringing out the arithmetic nature of the problem. Although these sums seem extremely hard to be computed individually, we manage to conclude by evaluating an average corresponding to the trace of finite intertwining.
@@ 1.5. Intertwining and an exponential sum
Assume that $p$ odd and $E/\mathbf{Q}_{p}$ unramified. Let $\lambda$ be a character of $E^{\times}$ sending $p$ to 1-1 and whose restriction to $\mathcal{O}^{\times}$ equals $\chi_{E}$, where $\chi:\mathcal{O}^{1}\twoheadrightarrow\mathbf{F}_{p^{2}}^{1}\to\mathbf{C}^{\times}$ is a (non-trivial) tamely ramified character. Letting $\pi_{n}$ denote the non-tempered representation of the Arthur packet $\Pi(\lambda)$, one has $\dim\pi_{n}^{K_{T}}=1$ where $K_{T}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&p\mathcal{O}&p\mathcal{O}\\ p\mathcal{O}&\mathcal{O}^{\times}&p\mathcal{O}\\ p\mathcal{O}&p\mathcal{O}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$.
@@
One has
$B\backslash G/K_{T}=\overline{B}\backslash\overline{G}/\overline{T}=\left\{\mathbf{1},w,\overline{[0,1]}=\left(\begin{smallmatrix}1&0&0\\ 0&1&0\\ \xi&0&1\end{smallmatrix}\right),\sigma_{y}=\overline{[1,y]}=\left(\begin{smallmatrix}1&0&0\\ -1&1&0\\ \xi y-\frac{1}{2}&1&1\end{smallmatrix}\right)\Big|y\in\mathbf{F}_{p}\right\}.$
Moreover, for any non-trivial $\chi$, the $K_{T}$-invariants in $\operatorname{Ind}_{B}^{G}(\mu)$ are supported by the double cosets of $\{\sigma_{y},y\in\mathbf{F}_{p}\}$ and, in addition if $\chi$ is quadratic, by the double coset of $\overline{[0,1]}$.
@@
Mackey’s Theorem and Frobenius Reciprocity yield that the dimension of $K_{T}$-invariants in $\operatorname{Ind}_{B}^{G}(\mu)$ equals the number of cosets $\sigma]\in B\backslash G/K_{T}$ such that $\mu$ has trivial restriction to $B\cap\sigma K_{T}\sigma^{-1}\supset Z$. As $\mu$ is ramified this is never the case for $\sigma=\mathbf{1}$, nor for $\sigma=w$, while $\sigma=\left(\begin{smallmatrix}1&0&0\\ 0&1&0\\ \xi&0&1\end{smallmatrix}\right)$ works if and only if $\chi$ is quadratic.
If $b=\left(\begin{smallmatrix}\bar{\alpha}&*&*\\ 0&1&*\\ 0&0&\alpha^{-1}\end{smallmatrix}\right)\in\sigma K_{T}\sigma^{-1}$, for some $y\in\mathbf{F}_{p}$, performing a matrix multiplication shows that
$\sigma_{y}^{-1}b\sigma_{y}\equiv\begin{pmatrix}*&*&*\\ (\bar{\alpha}-1)&*&*\\ ?&(\alpha^{-1}-1)&*\end{pmatrix}\mod{p}\in K_{T},$
hence $\alpha\equiv 1\pmod{p}$, ensuring the triviality of $\mu$ on $B\cap\sigma_{y}K_{T}\sigma_{y}^{-1}$.
In summary, the dimension of $\left(\operatorname{Ind}_{B}^{G}(\mu)\right)^{K_{T}}$ is $p+1$ for $\chi$ quadratic, and $p$ otherwise. ∎
@@
Recall that $\pi_{n}$ is the Langlands quotient of $\operatorname{Ind}_{B}^{G}(\mu)$ whose other Jordan-Hölder constituent is $\pi_{2}$. As taking invariants by an open compact subgroup is an exact functor in the category of admissible representations, Lemma 1.11 implies that
$\dim\pi_{n}^{K_{T}}+\dim\pi_{2}^{K_{T}}=p(+1).$
To show that $\pi_{n}^{K_{T}}\neq\{0\}$ one could try computing $\dim\pi_{2}^{K_{T}}$ as $\pi_{2}$ is a generic discrete series. Unfortunately Miyauchi’s theory [19, 20] of conductors for $\operatorname{U}(3)$-representations with respect to the paramodular groups $K_{r}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&p^{-r}\mathcal{O}\\ p^{r}\mathcal{O}&\mathcal{O}^{\times}&\mathcal{O}\\ p^{r}\mathcal{O}&p^{r}\mathcal{O}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$ predicts that the ramified $\pi_{n}$ and $\pi_{c}$ have no level, i.e. they have no invariants by $K_{r}$ for any $r$, while the level of $\pi_{2}$ is given by its conductor. When $\chi$ is the quadratic character, the $L$-parameter $(\lambda\otimes\mathrm{St})\oplus\mathbf{1}$ has conductor $2$, therefore the generic member $\pi_{2}$ in this $L$-packet has one dimensional invariants by $K_{2}$, hence also by $\gamma^{-1}K_{2}\gamma=\left(\begin{smallmatrix}\mathcal{O}^{\times}&p\mathcal{O}&\mathcal{O}\\ p\mathcal{O}&\mathcal{O}^{\times}&p\mathcal{O}\\ \mathcal{O}&p\mathcal{O}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G\supset K_{T}$. For other ramified, $\chi$’s $\pi_{2}$ has an invariant line by $K_{3}$ which has same volume as $K_{T}$ but is not conjugated to it, thus non-settling the non-vanishing of $\pi_{2}^{K_{T}}$ let alone computing its dimension.
Recall the Jacquet functor given by
$\operatorname{Ind}_{B}^{G}(\mu)\longrightarrow\mathbf{C}\cdot\mu\oplus\mathbf{C}\cdot\mu^{w},\quad f\mapsto(f(1),(Mf)(1))$
where the standard intertwining operator $M:\operatorname{Ind}_{B}^{G}(\mu)\to\operatorname{Ind}_{B}^{G}(\mu^{w})$ is defined via analytic continuation, as follows. For $s\in\mathbf{C}$, letting $\mu_{s}=\mu\cdot\delta^{s/2}$, the intertwining operator
$M_{s}:\operatorname{Ind}_{B}^{G}(\mu_{s})\to\operatorname{Ind}_{B}^{G}(\mu_{s}^{w}),\quad f\mapsto\int_{N}f(wn\cdot)dn$
is absolutely convergent for $\Re(s)\gg 0$ and $G$-equivariant. Moreover, for any section $f_{s}\in\operatorname{Ind}_{B}^{G}(\mu_{s})$ such that for all $g\in G$ the function $f_{s}(g)$ is analytic in $s\in\mathbf{C}$, the function $(M_{s}(f_{s}))(g)$, a priori only defined for $\Re(s)\gg 0$, is a rational function in $p^{-s}$, hence extends to a meromorphic function on all of $\mathbf{C}$ with only possibly a finite number of poles independent of $f$ and of $g$. In fact, it continues as an intertwining operator, i.e.
@@
We refer to [1, §1] for more detail and proofs, and we will only use that

### v2

@@ 1.4. Higher Iwahori invariants via matrix coefficients
In this subsection, we assume that $E/\mathbf{Q}_{p}$ is unramified and we let $\lambda_{0}$ denote the unique quadratic unramified character of $E^{\times}$. By Iwasawa decomposition, the corresponding $\pi=\operatorname{Ind}_{B}^{G}(\mu_{0})$ has one dimensional invariants by any given maximal open compact subgroup $K$ of $G$. The following proposition states, depending on $K$, whether the $K$-invariant line belongs its sub-representation $\pi_{2}$ or maps non-trivially to the quotient $\pi_{n}$ (see §1.2 for notation). We recall that $K^{\circ}$ and $K^{\prime}$ are the two standard maximal compact subgroups, $K^{\circ}$ being the hyperspecial one, and that the standard Iwahori subgroup $I$ equals $K^{\circ}\cap K^{\prime}$.
@@
One has $\dim\pi_{2}^{K^{\prime}}=\dim\pi_{n}^{K^{\circ}}=1$.
@@
Applying the Jacquet functor to the exact sequence (2) allows one to see that both $\pi_{2}$ and $\pi_{n}$ have $1$-dimensional II-invariants. As by Cartan decomposition $G$ is generated by $K$ and $\gamma$, hence by $K^{\circ}$ and $K^{\prime}$, it follows that exactly one amongst $\pi_{2}^{I}$ and $\pi_{n}^{I}$ is fixed by $K^{\circ}$, while the other one is fixed by $K^{\prime}$. By Iwasawa decomposition, the restriction to $K$ yields an isomorphism $\pi\xrightarrow{\sim}\operatorname{Ind}_{B\cap K}^{K}(\mu_{0})$, under which a basis $f_{K}$ of $\pi^{K}$ is mapped to $\mathbf{1}_{K}$. Moreover, the line $(\pi^{\vee})^{K}$ admits a basis $f_{K}^{\vee}$ characterized by
$\langle f,f_{K}^{\vee}\rangle=\frac{1}{\sqrt{\operatorname{vol}(K)}}\int\limits_{K}f(k)dk.$
@@
$\sqrt{\operatorname{vol}(K\gamma^{n}K)}\langle\gamma^{n}\cdot f_{K},f_{K}^{\vee}\rangle=\sqrt{[K:(K\cap\gamma^{-n}K\gamma^{n})]}\int\limits_{K}f_{K}(k\gamma^{n})dk.$
Using Iwahori decomposition for all $n\geqslant 1$ we have $K:(K\cap\gamma^{-n}K\gamma^{n})]/[K:I]=p^{4n-3}$ (resp. $p^{4n-1}$), where $K=K^{\circ}$ (resp. $K^{\prime}$). Since $(\mu_{0}\delta^{1/2})(\gamma)=-p^{3}$ we have $f_{K}\in\pi_{2}$ if, and only if,
(5) $(\Phi^{K}_{n})_{n}\in\operatorname{L}^{2}(\mathbf{Z}_{\geqslant 0})\text{, where }\Phi^{K}_{n}=p^{5n}\cdot\int\limits_{K}f_{K}(\gamma^{-n}k\gamma^{n})dk.$
The proof of the Proposition is then completed by the following Lemma. ∎
@@
The last row of an element $k\in K^{\prime}$ is given by $(0,0,1)\cdot k=(p\cdot c_{1}(k),p\cdot c_{2}(k),c_{3}(k))$ with $c_{2}(k)\in\mathcal{O}$ and $(c_{1}(k),c_{3}(k))\in(\mathcal{O}\times\mathcal{O}){\smallsetminus}(\mathcal{P}\times\mathcal{P})$. For $j\geqslant 0$ we let
$K^{\prime}_{j}=\left\{k\in K^{\prime}\Big|c_{1}(k)\in\mathcal{P}^{j}\right\}\text{ and }K^{\prime\times}_{j}=K^{\prime}_{j}{\smallsetminus}K^{\prime}_{j+1}.$
@@
First, we compute the volume of $K^{\prime}_{j}$, for $j\geqslant 1$. Using Iwahori decomposition one finds that
@@
Therefore $\frac{1}{\operatorname{vol}(K^{\prime})}\int\limits_{K^{\prime}}f_{K}(\gamma^{-n}k\gamma^{n})dk=\frac{1}{[K^{\prime}:K^{\prime}_{2n}]}+p^{-6n}\sum_{j=0}^{2n-1}(-1)^{j}p^{3j}\frac{1}{[K^{\prime}:K^{\prime\times}_{j}]}=$
@@
Alternatively, one could use MacDonald’s formula for zonal spherical functions to see that $\pi_{2}$ does not admit non-zero vectors fixed by the hyperspecial maximal open compact subgroup $K^{\circ}$. Indeed, [16, §5.5] allows one to write $\Phi^{K^{\circ}}_{n}$ from (5), up to a non-zero constant, as
$\Gamma_{\mu_{0}}\cdot\mu_{0}(\gamma^{-n})+\Gamma_{\mu_{0}^{\mathsf{w}}}\cdot\mu_{0}^{\mathsf{w}}(\gamma^{-n}),\text{ with }\Gamma_{\nu}=\frac{1-p^{2}\cdot\nu(\gamma^{-2})}{1-\nu(\gamma^{-2})}\cdot\frac{1-p^{2}\cdot\nu(\gamma^{-1})}{1-\nu(\gamma^{-1})},$
where the two factors in $\Gamma_{\nu}$ correspond respectively to the positive roots $\zeta$ and $2\zeta$ of $G$ (see §1.1). As $\mu_{0}(\gamma^{-1})=\mu_{0}^{\mathsf{w}}(\gamma)=-p^{-1}$, one has $\Gamma_{\mu_{0}}=0\neq\Gamma_{\mu_{0}^{\mathsf{w}}}$, hence the sequence $(\Phi^{K^{\circ}}_{n})_{n\geqslant 0}$ is not $\operatorname{L}^{2}$. This also shows, in passing, that $\pi_{n}$ is not square integrable.
@@
For $r\geqslant 1$ and for $\pi_{n}\in\Pi(\lambda_{0})=\Pi(\lambda_{0},\mathbf{1})$, one has $\dim\left(\pi_{n}^{I_{2r}}\right)\geqslant r+1$.
@@
Let $f$ be a basis of $\pi_{n}^{K^{\circ}}$ (see Proposition 1.5). For all $r\in\mathbf{Z}$, $\pi_{n}$ contains a (unique) line fixed by $\gamma^{r}K^{\circ}\gamma^{-r}$ (having $\gamma^{r}\cdot f_{K^{\circ}}$ as basis) and moreover, the stabilizer in $G$ of that line is $\gamma^{r}K^{\circ}\gamma^{-r}$. We claim that the vectors $f,\gamma\cdot f,\dots,\gamma^{r}\cdot f\in\pi_{n}$ are linearly independent. Indeed, if $f$ was a linear combination of the remaining $r$ vectors then it would be fixed by $\bigcap_{1\leqslant j\leqslant r}\gamma^{i}K^{\circ}\gamma^{-i}$ which is not contained in $K^{\circ}$. As any of these $(r+1)$ vectors is fixed by $I_{2r}$, the claim follows.
A similar claim holds for $\pi_{2}$ and can be proven exactly as above, using $K^{\prime}$ instead of $K^{\circ}$. ∎
While the unramified $A$-packet $\Pi(\lambda_{0})$ would suffice for our global applications when $D\equiv 3\pmod{8}$, the case of discriminants $D\equiv 7\pmod{8}$ would require the use of certain tamely ramified $A$-packets $\Pi(\lambda)$ and providing explicit levels for them is the object of the next subsection.
1.5. Intertwinings and exponential sums
Our arithmetic applications will require to show existence of non-zero $I_{r}$-invariants, for some $r\in\mathbf{Z}_{\geqslant 1}$, in certain ramified $A$-packets. This is delicate because of the lack of a new-vector theory for non-generic representations $\pi$ (see Remark 1.11). Also Casselman’s result asserting that $\pi^{I_{r}}$ surjects onto $\pi_{N}^{T\cap I_{r}}$ is inconclusive here as the latter vanishes, in contrast with the situation in [11] where the open compact is a pro-$p$-Iwahori subgroup of a sufficiently deep level. We will instead employ explicit methods and prove in Theorem 1.9 that $\pi_{n}$ has non-zero invariants by $K_{T}$, which contains a conjugate of $I_{3}$. It is relatively straightforward to determine all such vectors in the full induced representation, but it becomes a thorny comptutation to find a non-square integrable matrix coefficient as in §1.4. By another result of Casselman, matrix coefficients can be expressed in terms of the corresponding ones in the Jacquet module, given here by an explicit scalar product on $\pi_{N}\simeq\mathbf{C}^{2}$. Making this actually work requires non-vanishing of the second coordinate of the image of a $K_{T}$-invariant vector under the Jacquet functor which, once verified, leads directly to the result we seek.
We reduce the computation of the Jacquet functor to a precise statement about the intertwining operator at the level of finite reductive groups. The latter involves showing non-vanishing of some explicit exponential sums, bringing out the arithmetic nature of the problem. Although these sums seem extremely hard to compute individually, we manage to deduce it by precisely computing a suitable average of such sums corresponding to the trace of the finite intertwining operator.
@@ 1.5. Intertwinings and exponential sums
Assume that $p$ odd and $E/\mathbf{Q}_{p}$ unramified. Fix a character $\chi:E^{1}\twoheadrightarrow\mathbf{F}^{1}_{p^{2}}\to\mathbf{C}^{\times}$ such that $\chi^{3}\neq\mathbf{1}$ and let $\lambda=\lambda_{0}\cdot\chi_{E}$. Letting $\pi_{n}$ denote the non-tempered representation of the Arthur packet $\Pi(\lambda)$, one has $\dim\pi_{n}^{K_{T}}=1$.
@@
Consider the character $\mu_{\lambda}$ from (1) attached to $\lambda$ and trivial on $E^{1}$. One has
$B\backslash G/K_{T}=\overline{B}\backslash\overline{G}/\overline{T}=\left\{\mathbf{1}_{3},\mathsf{w},\sigma_{\infty}=\left(\begin{smallmatrix}1&0&0\\ 0&1&0\\ \xi&0&1\end{smallmatrix}\right),\sigma_{y}=\left(\begin{smallmatrix}1&0&0\\ -1&1&0\\ \xi y-\frac{1}{2}&1&1\end{smallmatrix}\right)\Big|y\in\mathbf{F}_{p}\right\}.$
For any non-trivial $\chi$, the $K_{T}$-invariant functions in $\operatorname{Ind}_{B}^{G}(\mu_{\lambda})$ are supported by the double cosets of $\{\sigma_{y}|y\in\mathbf{F}_{p}\}$ and, in addition when $\chi$ is quadratic, by the double coset of $\sigma_{\infty}$.
@@
By the Iwasawa decomposition, restriction to $K$ yields an isomorphism $\operatorname{Ind}_{B}^{G}(\mu_{\lambda})\xrightarrow{\sim}\operatorname{Ind}_{B\cap K}^{K}(\mu_{\lambda})$. As $\mu_{\lambda|B\cap K}$ factors through a character $\overline{\mu}_{\lambda}$ of $\overline{B}$, the space of $K_{1}$-invariants in $\operatorname{Ind}_{B\cap K}^{K}(\mu_{\lambda})$ is naturally identified with $\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})$, and therefore its subspace of $K_{T}$-invariants is given by $\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})^{\overline{T}}$. By definition, a double coset $\overline{B}\sigma\overline{T}$ supports a non-zero function contributing to the latter space if, and only if, $\overline{\mu}_{\lambda}$ is trivial on $\overline{B}\cap\sigma\overline{T}\sigma^{-1}$ (this is in accordance with Mackey’s Theorem for finite groups). As $\overline{\mu}_{\lambda}$ is ramified this is never the case for $\sigma=\mathbf{1}_{3}$ nor for $\sigma=\mathsf{w}$, while a matrix computation shows that $\sigma=\sigma_{\infty}$ works if, and only if, $\chi$ is quadratic. For $y\in\mathbf{F}_{p}$ and $k\in K_{T}$, writting $(k\bmod{p})=\operatorname{diag}\left(\bar{\alpha},\beta,\alpha^{-1}\right)$, we find
$(\sigma_{y}k\sigma_{y}^{-1}\bmod{p})=\left(\begin{smallmatrix}\bar{\alpha}&0&0\\ \beta-\bar{\alpha}&\beta&0\\ *&\beta-\alpha^{-1}&\alpha^{-1}\end{smallmatrix}\right)$
ensuring the triviality of $\mu_{\lambda}$ on $B\cap\sigma_{y}K_{T}\sigma_{y}^{-1}$, as in forces $\bar{\alpha}\equiv\beta\pmod{p}$ on such elements. ∎
Recall that $\pi_{n}$ is the Langlands quotient of $\operatorname{Ind}_{B}^{G}(\mu_{\lambda})$ whose other Jordan-Hölder constituent is the discrete series $\pi_{2}$. As taking invariants by an open compact subgroup is an exact functor in the category of admissible representations, Lemma 1.10 implies that
(6) $\dim\pi_{n}^{K_{T}}+\dim\pi_{2}^{K_{T}}=\begin{cases} p&\text{ if } \chi^{2}\neq\mathbf{1},\\ p+1&\text{if } \chi\text{ quadratic}.\end{cases}$
@@
To show $\pi_{n}^{K_{T}}\neq\{0\}$ one could instead try computing $\dim\pi_{2}^{K_{T}}$. Miyauchi’s theory [20, 21] of conductors with respect to the paramodular groups $K^{\prime}_{r}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&p^{-r}\mathcal{O}\\ p^{r}\mathcal{O}&\mathcal{O}^{\times}&\mathcal{O}\\ p^{r}\mathcal{O}&p^{r}\mathcal{O}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$ asserts that the level of the generic $\pi_{2}$ is given by its conductor, while $\pi_{c}$ and the ramified $\pi_{n}$ have no level, i.e. they have no non-zero invariants by $K^{\prime}_{r}$ for any $r$. When $\chi$ is the quadratic character, the $L$-parameter $(\lambda\otimes\mathrm{St})\oplus\mathbf{1}$ has conductor $2$, therefore the generic member $\pi_{2}$ in this $L$-packet has $1$-dimensional invariants by $K^{\prime}_{2}$, hence also by $\gamma^{-1}K^{\prime}_{2}\gamma=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{P}&\mathcal{O}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{P}\\ \mathcal{O}&\mathcal{P}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G\supset K_{T}$. For other ramified $\chi$’s, $\pi_{2}$ has an invariant line by $K^{\prime}_{3}$, having the same volume as $K_{T}$ but not conjugate to it, thus not settling the non-vanishing of $\pi_{2}^{K_{T}}$, let alone computing its dimension.
In order to show that $\pi_{n}$ admits non-zero $K_{T}$-invariants, it would be enough to consider the image of a $K_{T}$-invariant vector from Lemma 1.10 under the Jacquet functor and show that its second coordinate does not vanish. As well known, this functor is given by
$\operatorname{Ind}_{B}^{G}(\mu_{\lambda})\longrightarrow\mathbf{C}\cdot\mu_{\lambda}\delta^{1/2}\oplus\mathbf{C}\cdot\mu_{\lambda}^{\mathsf{w}}\delta^{1/2},\quad f\mapsto(f(1),(Mf)(1))$
where the standard intertwining operator $M:\operatorname{Ind}_{B}^{G}(\mu_{\lambda})\to\operatorname{Ind}_{B}^{G}(\mu_{\lambda}^{\mathsf{w}})$ is defined via analytic continuation, as follows. For $s\in\mathbf{C}$, letting $\mu_{\lambda,s}=\mu_{\lambda}\cdot\delta^{s/2}$, the intertwining operator
$M_{s}:\operatorname{Ind}_{B}^{G}(\mu_{\lambda,s})\to\operatorname{Ind}_{B}^{G}(\mu_{\lambda,s}^{\mathsf{w}}),\quad f_{s}\mapsto\int\limits_{N}f_{s}(\mathsf{w}n\cdot)dn$
is absolutely convergent for $\Re(s)\gg 0$ and $G$-equivariant. Moreover, for any section $f_{s}\in\operatorname{Ind}_{B}^{G}(\mu_{\lambda,s})$ such that for all $g\in G$ the function $f_{s}(g)$ is analytic in $s\in\mathbf{C}$, the function $(M_{s}(f_{s}))(g)$, a priori only defined for $\Re(s)\gg 0$, is a rational function in $p^{-s}$, hence extends to a meromorphic function on all of $\mathbf{C}$ with only possibly a finite number of poles independent of $f$ and of $g$. In fact, it continues as an intertwining operator, i.e.
@@
We refer to [1, §1] for more detail and proofs. We will computing explicitly the right hand side of